how many ounces of a 35% alcohol solution must be mixed with 17 ounces of a 40% alcohol solution to make 36% alcohol solution?
I just a formula to calculate this please! Thank you for ur help!

Answer by stanbon(75887)

how many ounces of a 35% alcohol solution must be mixed with 17 ounces of a 40% alcohol solution to make 36% alcohol solution?
-----------------
Equation:
alcohol + alcohol = alcohol
0.40*17 + 0.35x = 0.36(17+x)
-----
Multiply thru by 100 to get:
40*17 + 35x = 36*17 + 36x
x = 4*17
---
x = 68 ounces (amt. of 35% solution needed for the mixture)
==============================================================
Cheers,
Stan H.